OpenStudy (anonymous):

help whats the function of y'= 2x^2-x

OpenStudy (anonymous):

y'=y'1+y'2 where y'1=2x^2 and y'2=-x y1=2/3x^3 y2=-x^2/2 y=2/3x^3-x^2/2

OpenStudy (anonymous):

ty